How to find the volume of HCl? and Borax? S? Ksp? Concentration Borax was not given but 14 grams was used in the solution mixed with 120 ml. The concentration of HCl is 0.5 M.

Warm Water

Temperature

60

Trial 1

Trial 2

Trial 3

Trial 4

Volume of Borax

Final burette reading

7.98

7.81

7.99

7.89

Initial burette reading

0.03

0.02

0.02

0.01

Volume HCl

S = [B4O5(OH)4]2-

Mean [B4O5(OH)4]2-

Std Deviation

KSp
Answer #1

* Volume of HCl for trial 1(V2) = final - initial burette reading = 7.98 - 0.03 =7.95 ml

* Volume of HCl for trial 2 (V4) = 7.81 - 0.02 = 7.78 ml

* Volume of HCl for trial 3 (V6) = 7.99 - 0.02 = 7.97 ml

* Volume of HCl for trial 4 (V8) = 7.89 - 0.01 = 7.88 ml

To get the volume of borax , 1st we have to find the concentration of borax And the concentration will be in terms of Molarity (M) because the concentration of hcl is also given in M.

Amount of borax = molecular mass/ 1000 × molarity × 120 ml

14g = 381.37 g/1000 L × molarity × 120ml

Molarity = 14 / 0.38 × 120

Molarity = 14/ 45.6 = 0.3M

Therefore the concentration of borax is 0.3M.

Volume of Borax can be calculated with molarity equation [M1V1=M2V2].

* Volume of Borax in trial 1(V1)

M1V1 = M2V2 [ M1V1 is the molarity and volume of Borax while M2V2 is molarity and volume of HCl]

0.3 × V1 = 0.5 × 7.95

V1 = 3.98/0.3 = 13.25 ml

* Volume of borax in trial 2 (V3)

M1V3 = M2V4   

0.3 × V3 = 0.5 × 7.79

V3 = 3.895/ 0.3 = 12.98 ml

* Volume of borax in trial 3 (V5)

M1V5 = M2V6

0.3 × V5= 0.5 × 7.97

V5 = 3.985/ 0.3 = 13.28 ml

* Volume of borax in trail 4 (V7)

M1V7 = M2V8

0.3 × V7 = 0.5 × 7.88

V7 = 3.94/ 0.3 = 13.13 ml

solubility (s) is define as the concentration of substance in a saturated solution . It is generally represented as grams of solute in moles per litre i.e, molarity .. therefore the solubility of borax is 0.3M .

And the solubility product (ksp) can be calculated as follows :

Na2[B4O5(OH)4].8H2O ------> 2Na+ + [B4O5(OH)4]2-

here 2Na+ = 2s and [B4O5(OH)]2- = s (were s is solubility )

Ksp = [Na]2 [B4O5(OH)4]

Ksp=s(2s)2

Ksp= 4s3= 4 × (0.3)3 = 4 × 0.027= 0.108